can you help me with this problem please.
x^2 + 12x + 36 = 0
thank you
julie and have a happy new year

Answer by stanbon(75887)

x^2 + 12x + 36 = 0
Factor:
(x+6)(x+6)= 0
since the product is zero one of the factors must be zero.
x+6=0
x = -6 (twice)
==================
Happy New Year!!!
===================
Cheers,
Stan H.